OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (jhannybean):

hint: \(\large \sin(3x)=\sin(2x+x)=\sin(2x)\cos(x)+\cos(2x)\sin(x)\)

OpenStudy (anonymous):

Could you please explain further? I have a difficult time grasping trig concepts.

OpenStudy (jhannybean):

Well, ok. you have \[\large \sin(3x)-\cos(x)\] and you can rewrite sin(3x) as the following, \[\large \sin(2x+x) = \sin(2x)\cos(x)+\cos(2x)\sin(x)\] So \[\large \sin(3x)-\cos(x) = \sin(2x)\cos(x)+\cos(2x)\sin(x)-\cos(x)\]

OpenStudy (anonymous):

So would it be 3 sin x cos x - sin^3x - cos x?

OpenStudy (jhannybean):

Oh...hm..\[\large \sin(2x)\cos(x)+\cos(2x)\sin(x)-\cos(x)\]\[\large 2\sin(x)\cos(x)\cos(x)+(2\cos^{2}x-1)(\sin(x))-\cos(x)\]\[\large 2\sin(x)\cos^2(x)+2\cos^2(x)\sin(x)-\sin(x)-\cos(x)\]

OpenStudy (jhannybean):

Right? I'm going to need a little help here too... haha.

OpenStudy (anonymous):

I'm not sure. If it helps, the choices are: 2 sin x - sin3x - cos x 2 sin x cos2x + sin x - 2 sin3x - cos x 2 sin3x cos4x + 1 3 sin x cos x - sin3x - cos x

OpenStudy (jhannybean):

Bleh.

OpenStudy (anonymous):

Yeah, I'm really awful at these things. Thank you so much for your help, though!

OpenStudy (jhannybean):

I'll figure it out soon enough, lol.

OpenStudy (jhannybean):

bbs. dont leave the problem!

OpenStudy (anonymous):

So, out of the choices I listed about, which would it be? Thanks for your help!

OpenStudy (zzr0ck3r):

o there are choices lol , one sec

OpenStudy (jhannybean):

hahaha. wooow. good job zz

OpenStudy (zzr0ck3r):

um im confused, all the answers are not in terms of only cos(x) sin(x)

OpenStudy (anonymous):

How are they not?

OpenStudy (zzr0ck3r):

sin(3x) in the first one cos(2x) in the second sin(3x) in the third sin(3x) in the fourth

OpenStudy (zzr0ck3r):

im confused I think... is this sin(3x) or sin(x)^3?

OpenStudy (anonymous):

Choice A: 2 sin x - sin^3x - cos x

OpenStudy (anonymous):

Choice B: 2 sin x cos^2x + sin x - 2 sin^3x - cos x

OpenStudy (unklerhaukus):

\[\sin(3x)-\cos(x)\\ =\sin(2x)\cos(x)+\cos(2x)\sin(x)-\cos(x)\\ =\Big(\sin(x)\cos(x)+\cos(x)\sin(x)\Big)\cos(x)+\Big(1-2\sin^2(x)\Big)\sin(x)-\cos(x)\\ =2\sin(x)\cos^2(x)+\sin(x)-2\sin^3(x)-\cos(x)\]

OpenStudy (anonymous):

Thank you so much, UnkleRhaukus!

OpenStudy (anonymous):

It's okay!!

OpenStudy (unklerhaukus):

i have used\[\sin(a+b)=\sin(a)\cos(b)+\cos(a)\sin(b)\] like @Jhannybean suggested and\[\cos(2u)=1-2\sin^2(u)\]

OpenStudy (jhannybean):

such a hard question!!